OpenStudy (angel_kitty12):

Can someone help ASAP I don't understand this problem: 5 sqrt x^5/2y^3

OpenStudy (angel_kitty12):

How the problem looks

OpenStudy (campbell_st):

so what are you being asked to do..?

OpenStudy (angel_kitty12):

The topic is nth roots and rational exponents

Directrix (directrix):

@Angel_Kitty12 What are the specific instructions for this posted problem? Thanks.

OpenStudy (campbell_st):

ok...s o you can simplify this to \[\frac{\sqrt[5]{x^5}}{\sqrt[5]{2y^3}}\] does that make sense...

OpenStudy (campbell_st):

if you want to write rewrite the denominator you can use fractional indices \[\frac{\sqrt[5]{x^5}}{(2y^3)^{\frac{1}{5}}}\]

OpenStudy (campbell_st):

simplify the numerator and re-write the denonimator with a negative index